OpenStudy (anonymous):

How do I express F(X)=x^2+3x-2 if the given domain is F(x-1)?

OpenStudy (anonymous):

Do Sumthing by ur own!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (hoblos):

solved before

OpenStudy (anonymous):

holblos let her do sumthing by her own.....she want to get everything done through others

OpenStudy (hoblos):

i agree with u sheg no more answers Ms. taylor we will just tell u the way

OpenStudy (anonymous):

hey u will not even get a single answer if we want but we want that u sud try on ur own so that u will learn something it is for you not for us..........so take this optimistically

OpenStudy (anonymous):

I only asked for answers to the one's that I am weak in.

OpenStudy (anonymous):

just replace the value of x like in this case x = x-1

OpenStudy (anonymous):

Kay

OpenStudy (anonymous):

F(X)=x^2+3x-2 \[F(x-1) = (x-1)^2 + 3(x-1) - 2\]\[=> F(x-1) = x^2 -2x +1 +3x -3 -2\]\[=> F(x-1) = x^2 + x -4\]